OpenStudy (anonymous):

Find dy/dx by implicit differentiation. 3cos(8x)sin(4y)=3

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

Differentiate both sides and solve for y':\begin{eqnarray*}&&\left(3\cos(8x)\sin(4y)\right)'=(3)' \\ \Rightarrow &&3\left(\left(\cos(8x)\right)'\sin(4y) + \cos(8x)\left(\sin(4y)\right)' \right) = 0\\\Rightarrow&&3\left( -8\sin(8x)\sin(4y) + 4\cos(8x)\cos(4y)y'\right) = 0\\\Rightarrow&&y' = 2\tan(8x)\tan(4y).\end{eqnarray*}

OpenStudy (anonymous):

\[24\sin(8x)\sin(4y) + 12\cos(8x)\cos(4y)y'=0\] solve for \[y'\]

OpenStudy (anonymous):

krebante is right. i made a mistake the the first term should be \[-24\sin(8x)\sin(4y)\] sorry.

OpenStudy (anonymous):

hey thanks:) can u guys help me on another problem